OpenStudy (anonymous):

Shereda bought several dozen roses and several dozen carnations. The roses cost $15 per dozen, and the carnations cost $8 per dozen. Shereda bought a total of 17 dozen flowers and paid a total of $192. How many roses did she buy?

OpenStudy (anonymous):

Write down what you know, and assign letters to your variables: r = number of roses (in dozens) c = number of carnations (in dozens) Total cost = T = 192 Total number of flowers (in dozens) = F = 17 Then: 15r + 8c = 192 15r = 192 - 8c 15r/8 = 24 - c c = 24 - 15r/8 r + c = 17 c = 17 - r Therefore: 17 - r = 24 - 15r/8 -r + 15r/8 = 24 - 17 7r/8 = 7 r = 8 dozen roses

OpenStudy (anonymous):

why is it 7r/8 at the end?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Because -r + 15r/8 = 15r/8 - 8r/8 = 7r/8